OpenStudy (anonymous):

what polynomial has roots of -4, 1, and 6? a: x^3-3x^2-22x+24 b: x^3 - x^2 - 26x - 24 c: x^3 + x^2 - 26x + 24 d: x^3 +3x^2 + 14x - 24

jimthompson5910 (jim_thompson5910):

you could use the zero product property in reverse, then expand but since you're given the answer choices you can just check each one

jimthompson5910 (jim_thompson5910):

all you have to do is plug each root into each equation if you get 0 for each root on a particular equation, then you have found your answer

OpenStudy (anonymous):

plug each root in for x?

jimthompson5910 (jim_thompson5910):

for instance, in choice A, the equation is y = x^3-3x^2-22x+24 if -4 was a root, then plugging x = -4 into the equation *should* give you y = 0 if it doesn't give you y = 0, then x = -4 isn't a root

OpenStudy (espex):

As @whpalmer4 said, your expression is equal to \((x-r_1)(x-r_2)(x-r_3)...(x-r_k)\) so you can plug in your roots and expand out to get your equation.

jimthompson5910 (jim_thompson5910):

you have to do this with all 3 roots (for each answer choice)

OpenStudy (anonymous):

im confused between reading his then yours

jimthompson5910 (jim_thompson5910):

go with whichever way is easiest for you

OpenStudy (whpalmer4):

Two different approaches. Both are valid. One of is practical only because you have a list of answer choices that you can try.

OpenStudy (anonymous):

so is B correct?

jimthompson5910 (jim_thompson5910):

what do you get when you plug in x = 1 into choice B?

OpenStudy (anonymous):

no zero haha. -50 :o

jimthompson5910 (jim_thompson5910):

so x = 1 is NOT a root of x^3 - x^2 - 26x - 24 which means you can eliminate B

jimthompson5910 (jim_thompson5910):

see how that works?

jimthompson5910 (jim_thompson5910):

the result has to be zero if that x value is a root

OpenStudy (anonymous):

well had C to start out with and you said its not A

jimthompson5910 (jim_thompson5910):

when did I say it's not A? I never said either way

OpenStudy (anonymous):

well its not A because -4 doesnt work

OpenStudy (anonymous):

Plug in roots into the equations; only one will consistently yield 0 for each root.

jimthompson5910 (jim_thompson5910):

what do you get when you plug x = -4 into choice A

OpenStudy (anonymous):

-50

OpenStudy (anonymous):

im obviously doing it wrong because -4 doesnt make any of them zero for me

jimthompson5910 (jim_thompson5910):

how are you typing it into your calculator

OpenStudy (anonymous):

im writing it out on paper

jimthompson5910 (jim_thompson5910):

is it possible to show us what you're writing down so we can see where you're going wrong

OpenStudy (anonymous):

ok your not helping me. I knew how to do the equation I came on here because like I just said im doing something rong. Im plugging -4 in for the X there isnt that much more to explain

jimthompson5910 (jim_thompson5910):

ok sry

jimthompson5910 (jim_thompson5910):

just keep in mind that x^3 turns into (-4)^3 = -64 and something like x^2 turns into (-4)^2 = 16

jimthompson5910 (jim_thompson5910):

please try not to be so rude next time

OpenStudy (anonymous):

ive narrowed it down to c and D. and Im sorry If I came off As rude but you werent helping me,

OpenStudy (anonymous):

is it C @eSpeX

OpenStudy (espex):

I did not get C,.

OpenStudy (espex):

I did, \((x-(-4))(x-1)(x-6)\)

OpenStudy (anonymous):

then it must be D because A or B didnt work for me

OpenStudy (espex):

When you foil this, what do you get?

OpenStudy (espex):

\((x^2+3x-18x)(x-6)\)

OpenStudy (espex):

oops, I meant, (x^2+3x-4)(x-6)

OpenStudy (whpalmer4):

Let's try evaluating A again. You say it doesn't work for -4 for you? \[(-4)^3-3(-4)^2-22(-4)+24=-64-3(16)-22(-4)+24=\]

OpenStudy (espex):

Okay, work through with @whpalmer4 and if you still haven't an answer we can try again. :)

OpenStudy (anonymous):

se Thats where I messed up I just got 0 for A for -4

OpenStudy (whpalmer4):

You should get 0 if it is a root...

OpenStudy (whpalmer4):

Try it with the other two roots as well. If you get 0 for them as we'll, then A would be your answer. The value of the polynomial is 0 at all roots.

OpenStudy (anonymous):

for 1 in A I got 12? is that incorrect?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Afraid so...

OpenStudy (whpalmer4):

\[(1)^3-3(1)^2-22(1)+24=1-3-22+24\]

OpenStudy (anonymous):

So its A?

OpenStudy (whpalmer4):

Only if you try all 3 and get 0 for each...

OpenStudy (whpalmer4):

Do you agree that 1 gives you 0 in A?

OpenStudy (anonymous):

Yes I solved it

OpenStudy (whpalmer4):

Okay, how about 6?

OpenStudy (whpalmer4):

\[(6)^3-3(6)^2-22(6)+24=216-108-132+24=\]

OpenStudy (anonymous):

Yup Its A!

OpenStudy (whpalmer4):

:-)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (whpalmer4):

The other approach to this was to do polynomial multiplication, namely \[(x+4)(x-1)(x-6)=(x^2-x+4x-4)(x-6)\]\[=.x^3-6x^2-x^2+6x+4x^2-24x-4x+24\]\[=x^3-3x^2-22x+24 \]

OpenStudy (whpalmer4):

Both offer significant opportunities for error if your algebra is mistake-prone :-)